Determine the area under the standard normal curve that lies to the right of
(a) z= -0.77 (b) z= -1.17 (c) z= -1.08 (d) z= -0.47
(a) The area to the right of z= -0.77 is ___ (Round to four decimal points as needed)
(b) The area to the right of z= -1.17 is ___ (Round to four decimal points as needed)
(c) The area to the right of z= -1.08 is ___ (Round to four decimal points as needed)
(d) The area to the right of z= -0.47 is ___ (Round to four decimal points as needed)

Answer by Edwin McCravy(20086)

Determine the area under the standard normal curve that lies to the right of
(a) z= -0.77 (b) z= -1.17 (c) z= -1.08 (d) z= -0.47
On your TI-84

Press CLEAR
Press 2nd
Press VARS
Press 2  

You should see this on your screen:

normalcdf(

after that type -0.77,99999999)

[Careful: be sure you type the key that has (-),
the 'negative' key, and NOT the 'minus' key above the +
key.  That (-) key is at the very bottom of the keyboard, 
just right of the "." key.]   

You should see this on your screen:

normalcdf(-0.77,
99999999)

Press ENTER

You should see this on your screen:

normalcdf(-0.77,
99999999)
     .7793501242

Round this off to .7794

(a) The area to the right of z= -0.77 is .7794


----------------------------------
Do the same for the others:

(b) The area to the right of z= -1.17 is .8790

(c) The area to the right of z= -1.08 is .8599 

(d) The area to the right of z= -0.47 is .6808 

Edwin